# Recursive Gaussian Process State Space Model

### v2

@@ Abstract
Learning dynamical models from data is not only fundamental but also holds great promise for advancing principle discovery, time-series prediction, and controller design. Among various approaches, Gaussian Process State-Space Models (GPSSMs) have recently gained significant attention due to their combination of flexibility and interpretability. However, for online learning, the field lacks an efficient method suitable for scenarios where prior information regarding data distribution and model function is limited. To address this issue, this paper proposes a recursive GPSSM method with adaptive capabilities for both operating domains and Gaussian process (GP) hyperparameters. Specifically, we first utilize first-order linearization to derive a Bayesian update equation for the joint distribution between the system state and the GP model, enabling closed-form and domain-independent learning. Second, an online selection algorithm for inducing points is developed based on informative criteria to achieve lightweight learning. Third, to support online hyperparameter optimization, we recover historical measurement information from the current filtering distribution. Comprehensive evaluations on both synthetic and real-world datasets demonstrate the superior accuracy, computational efficiency, and adaptability of our method compared to state-of-the-art online GPSSM techniques.
@@
$\bm{K}_{ab}$  = kernel matrix indexed by function outputs or function inputs
$\bm{k}_{ab}$  = $\bm{K}_{ab}\bm{K}_{bb}^{-1}$
$\bm{\theta}$  = kernel hyperparameters
$\bm{\Sigma}_{p}$  = process noise covariance
$\bm{\Sigma}_{m}$  = measurement noise covariance
$\bm{x}_{t}$  = system state at time $t$
$\bm{y}_{t}$  = measurement at time $t$
@@
$\bm{f}$  = all function values of $f(\cdot)$
$\bm{u}$  = inducing point set
@@
$\bm{\mu}_{t}$  = mean of the system state $\bm{x}$ at time $t$
$\bm{P}_{t}$  = covariance of the system state $\bm{x}$ at time $t$
$\bm{m}_{u}$  = mean of the inducing points $\bm{u}$
$\bm{V}_{xu}$  = cross-covariance between the system state $\bm{x}$ and the inducing points $\bm{u}$
$\bm{S}_{uu}$  = auto-covariance of the inducing points $\bm{u}$
$f_{t}$  = function value $f(\bm{\mu}_{t})$
$m_{f_{t}}$  = mean of the function value $f_{t}$
$\bm{F}_{t}$  = predicted mean of the state transition function, i.e., $\bm{F}_{t}=\bm{F}(\bm{\mu}_{t},m_{f_{t}})$
$\bm{A}_{f}$  = transition Jacobian with respect to the function value $f_{t}$
$\bm{A}_{x}$  = transition Jacobian with respect to the state $\bm{x}$
$\bm{C}_{x}$  = measurement Jacobian with respect to the state $\bm{x}$
$\bar{\bm{u}}$  = augmented inducing point set, i.e., $\bar{\bm{u}}=[\bm{u}^{T},f_{t}^{T}]^{T}$
$\bm{X}_{t}$  = augmented state at time $t$, i.e., $\bm{X}_{t}=[\bm{x}_{t}^{T},{\bm{u}}^{T}]^{T}$
$\bm{\xi}_{t}$  = mean of the augmented state $\bm{X}_{t}$
$\bm{\Sigma}_{t}$  = covariance of the augmented state $\bm{X}_{t}$
$\bar{\bm{X}}_{t}$  = augmented state with $\bar{\bm{u}}$, i.e., $\bar{\bm{X}}_{t}=[\bm{x}_{t}^{T},\bar{\bm{u}}^{T}]^{T}$
$\bar{\bm{\xi}}_{t}$  = mean of the augmented state $\bar{\bm{X}}_{t}$
$\bar{\bm{\Sigma}}_{t}$  = covariance of the augmented state $\bar{\bm{X}}_{t}$
$\bm{\Phi}$  = transition Jacobian for the augmented state $\bar{\bm{X}}_{t}$
$\bm{\Sigma}_{p,\bm{X}}$  = process noise covariance for the augmented state $\bar{\bm{X}}_{t}$
$\bm{H}$  = measurement Jacobian for the augmented state $\bar{\bm{X}}_{t}$
$M$  = budget for inducing point set size
$u_{\mathrm{d}}$  = inducing point to be discarded
$\bm{u}_{\mathrm{l}}$  = inducing points left after discarding
@@
$s_{d}$  = score evaluting the importance of the inducing points
$\bm{Q}$  = inversion of the kernel matrix $\bm{K}_{uu}^{-1}$
$\bm{\Omega}$  = inversion of the joint covariance $\bm{\Sigma}_{t}^{-1}$
$\bm{Q}_{dd}$  = the $i_{\mathrm{d}}$-th diagonal element of the matrix $\bm{Q}$
$\bm{Q}_{du}$  = the $i_{\mathrm{d}}$-th row of the matrix $\bm{Q}$
$\bm{\Omega}_{dd}$  = the $(d_{x}+i_{\mathrm{d}})$-th diagonal element of the matrix $\bm{\Omega}$
$\gamma$  = novelty metric for inducing points
$\varepsilon_{\text{tol}}$  = threshold for adding inducing points
@@ 1 Introduction
[…]
State-space models (SSMs), also known as hidden Markov models (HMMs), describe the evolution and observation processes of latent states through a state transition model and a measurement model. By summarizing historical information into the latent state, SSMs provide a concise representation of system dynamics and are widely applied in fields such as aerospace [1], finance [2], neuroscience [3], target tracking [4], and control systems [5, 6]. Nevertheless, due to the complexity of real-world systems, establishing precise SSMs based on first principles is challenging. To achieve sufficiently accurate system models, extensive research has focused on data-driven SSM modeling [7, 8, 9].
@@
    Particle Filter (PF)-based method [38, 39, 40]. This approach is similar to the first, with the key difference being the use of the particle filter (PF) for the inference process. Since PF is inefficient at handling high-dimensional states, GP parameters are not augmented into the state space. Instead, each particle is assigned a GP, which is recursively learned using the particle state. This method can theoretically achieve precise inference for arbitrary observational distributions, provided the number of particles is sufficient. However, despite the inherent issues with PF, such as high computational cost and particle degeneracy [41], it also requires prior knowledge of the operating domain.
@@
To overcome the aforementioned shortcomings, a recursive learning algorithm for GPSSMs is proposed in this paper, capable of online adaptation to the operating domain and GP hyperparameters. The main contributions can be summarized as follows: Firstly, to implement a domain-independent recursive learning algorithm for GPSSMs, a novel Bayesian update equation is derived without pre-parameterizing the GP. In the derivation process, only mild linearization approximations are applied to achieve a closed-form solution, offering advantages in learning speed and stability. Through continuously expanding the inducing-point set, this method can accommodate real-time data distributions. Secondly, to address the computational cost issue caused by the increasing number of inducing points, an selection algorithm is developed with both addition and removal operations. To minimize accuracy loss, an optimal discarding equation and selection metric for inducing points are derived based on informative criteria. Thirdly, to alleviate the burden of offline hyperparameter tuning, an online GP hyperparameter optimization method is proposed. In this method, an approximate observation likelihood model is extracted from the obtained posterior distribution. Utilizing this likelihood model as the information source, the GP hyperparameters are optimized, and the posterior distribution can be adjusted as the hyperparameters change. Fourthly, to ensure numerical stability, a stable implementation method based on Cholesky decomposition is developed. Additionally, the extension to multi-output GP is provided, which significantly enhances the practicality and flexibility. Finally, we validate the effectiveness of the proposed method through two synthetic simulations and a real-world dataset. The experimental results demonstrate the superior performance of the proposed method in terms of learning accuracy, computational efficiency, and adaptive capacity for operating domains and hyperparameters.
@@ 3.2 Factorized Approximate Distribution
To address the nonlinearity within the transition and measurement model, first-order linearization is applied in this subsection, whose result is the following theorem.
@@
$q(\bm{x}_{t},\bm{f})=\int p(\bm{f}|\bm{u})q(\bm{x}_{t},\bm{u})\textbf{d}\bm{u}$ (7)
where $\bm{u}\triangleq\{u_{i}\}_{i=1}^{n_{u}}$ is a collection of a finite number of inducing points, $p(\bm{f}|\bm{u})$ is the exact GP conditional distribution similar to (2), namely,
$p(\bm{f}|\bm{u})=\mathcal{N}\left(\bm{K}_{fu}\bm{K}_{uu}^{-1}\bm{u},\,\bm{K}_{ff}-\bm{K}_{fu}\bm{K}_{uu}^{-1}\bm{K}_{uf}\right)$ (8)
@@
Given Theorem 1, since $p(\bm{f}|\bm{u})$ is known, we only need to update the finite-dimensional distribution $q(\bm{x}_{t},\bm{u})$, which build a foundation for addressing the non-parametric nature of the GP. Moreover, the approximate joint distribution remains Gaussian, which can be evaluated using closed-form update equations. In the following, we will prove Theorem 1 and then derive these equations.
@@
$q(\bm{x}_{t},\bm{u})$ $=\mathcal{N}\left(\begin{bmatrix}\bm{x}_{t}\\ \bm{u}\\ \end{bmatrix}\Bigg{|}\begin{bmatrix}\bm{\mu}_{t}\\ \bm{m}_{u}\\ \end{bmatrix},\begin{bmatrix}\bm{P}_{t}&\bm{V}_{xu}\\ \bm{V}_{xu}^{T}&\bm{S}_{uu}\\ \end{bmatrix}\right)$ (9)
$q(\bm{x}_{t},\bm{f})$ $=\mathcal{N}\left(\begin{bmatrix}\bm{x}_{t}\\ \bm{f}\\ \end{bmatrix}\Bigg{|}\begin{bmatrix}\bm{\mu}_{t}\\ \bm{m}_{f}\\ \end{bmatrix},\begin{bmatrix}\bm{P}_{t}&\bm{V}_{xf}\\ \bm{V}_{xf}^{T}&\bm{S}_{ff}\\ \end{bmatrix}\right)$
@@
By means of the above notation, we will illustrate the proof of Theorem 1, which is based on mathematical induction. First, it can be demonstrated that at time 00, the joint distribution naturally satisfies the factorized form (7):
@@
where we use the notation for moments in (9) and the shorthand notation $f_{t}=f(\bm{\mu}_{t})$, $m_{f_{t}}=\mathbb{E}_{q(\bm{x}_{t},\bm{u})}[f_{t}]$, $\bm{F}_{t}=\bm{F}(\bm{\mu}_{t},m_{f_{t}})$. Additionally, the two Jacobian matrices in the equation are
@@ 3.3 Moment Matching Equation
where $\bar{\bm{\xi}}_{t+1}^{-}$ and $\bar{\bm{\Sigma}}_{t+1}^{-}$ represent the predicted mean and covariance for the augmented state $\bar{\bm{X}}_{t+1}$. $\bm{\Phi}$ denotes the transition Jacobian matrix for $\bar{\bm{X}}_{t+1}$, $\bm{\Sigma}_{f,\bar{X}}$ is the process noise covariance for $\bar{\bm{X}}_{t+1}$, and $\bm{I}_{n_{u}}$ is .
@@ 4.1 Dynamic Adjusting the Inducing Points Set
$\hat{q}(\bm{x},\bm{f})=p(\bm{f}|\bm{u})\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})$ (19)
where $p(\bm{f}|\bm{u})$ is the conditional prior (8), and $\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})$ is defined to be Gaussian for achieving closed-form implementation. Then, to ensure minimal accuracy loss, we can seek the optimal solution within the distribution family of $\hat{q}(\bm{x},\bm{f})$, such that the distance between it and the original joint distribution $q(\bm{x},\bm{f})=p(\bm{f}|\bm{u})q(\bm{x},\bm{u})$ is minimized. Since $\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})$ controls the approximate distribution $\hat{q}(\bm{x},\bm{f})$, it is sufficient to find the optimal distribution for $\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})$. Here, we use the inclusive KL divergence as a distance metric, namely $\mathrm{KL}[q\|\hat{q}]=\int q\log\left(q/\hat{q}\right)$, which results in a simple optimal distribution:
@@
$=\arg\min\limits_{\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})}\mathrm{KL}\left[p(\bm{f}|\bm{u})q(\bm{x},\bm{u})\|p(\bm{f}|\bm{u}_{\mathrm{l}})\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})\right]$
@@ 4.2 Online Optimization of GP Hyperparameters
where $q_{\text{new}}(\bm{x}_{t},\bm{f})$ denotes the approximate joint distribution corresponding to the new hyperparameters $\bm{\theta}_{\text{new}}$. In (28), since the distributions $p(\bm{f};\bm{\theta}_{\text{old}})$, $p(\bm{f};\bm{\theta}_{\text{new}})$, and $q_{\text{old}}(\bm{x}_{t},\bm{f})$ are all Gaussian, the new approximate distribution $q_{\text{new}}(\bm{x}_{t},\bm{f})$ is also Gaussian. By expanding the distribution related to $\bm{f}$ in the third row of (28) with the conditional distribution $p(\bm{f}|\bm{u})$, it can be shown that the new distribution retains the factorized form $q_{\text{new}}(\bm{x}_{t},\bm{f})=p(\bm{f}|\bm{u};\bm{\theta}_{\text{new}})q_{\text{new}}(\bm{x}_{t},\bm{u})$, where:

### v3

@@ Abstract
Learning dynamical models from data is not only fundamental but also holds great promise for advancing principle discovery, time-series prediction, and controller design. Among various approaches, Gaussian Process State-Space Models (GPSSMs) have recently gained significant attention due to their combination of flexibility and interpretability. However, for online learning, the field lacks an efficient method suitable for scenarios where prior information regarding data distribution and model function is limited. To address this issue, this paper proposes a recursive GPSSM method with adaptive capabilities for both operating domains and Gaussian process (GP) hyperparameters. Specifically, first-order linearization is first applied to derive a Bayesian update equation for the joint distribution between the system state and the GP model, enabling closed-form and domain-independent learning. Second, an online selection algorithm for inducing points is developed based on informative criteria to achieve lightweight learning. Third, to support online hyperparameter optimization, historical measurement information is recovered from the current filtering distribution. Comprehensive evaluations on both synthetic and real-world datasets demonstrate the superior accuracy, computational efficiency, and adaptability of the proposed method compared to state-of-the-art online GPSSM techniques.
@@
$\bm{\theta}$  = kernel hyperparameters, nondimensional unit
$\bm{K}_{ab}$  = kernel matrix indexed by function outputs or function inputs, nondimensional unit
$\bm{k}_{ab}$  = $\bm{K}_{ab}\bm{K}_{bb}^{-1}$, nondimensional unit
$\bm{\Sigma}_{p}$  = process noise covariance, nondimensional unit
$\bm{\Sigma}_{m}$  = measurement noise covariance, nondimensional unit
$\bm{x}_{t}$  = system state at time $t$, nondimensional unit
$\bm{y}_{t}$  = measurement at time $t$, nondimensional unit
@@
$\bm{f}$  = all function values of $f(\cdot)$, nondimensional unit
$\bm{u}$  = inducing point set, nondimensional unit
$\bm{f}_{\neq u}$  = function values of $f(\cdot)$ excluding the inducing points $\bm{u}$, nondimensional unit
@@
$\bm{\mu}_{t}$  = mean of the system state $\bm{x}$ at time $t$, nondimensional unit
$\bm{P}_{t}$  = covariance of the system state $\bm{x}$ at time $t$, nondimensional unit
$\bm{m}_{u}$  = mean of the inducing points $\bm{u}$, nondimensional unit
$\bm{V}_{xu}$  = cross-covariance between the system state $\bm{x}$ and the inducing points $\bm{u}$, nondimensional unit
$\bm{S}_{uu}$  = auto-covariance of the inducing points $\bm{u}$, nondimensional unit
$f_{t}$  = function value $f(\bm{\mu}_{t})$, nondimensional unit
$m_{f_{t}}$  = mean of the function value $f_{t}$, nondimensional unit
$\bm{F}_{t}$  = predicted mean of the state transition function, i.e., $\bm{F}_{t}=\bm{F}(\bm{\mu}_{t},m_{f_{t}})$, nondimensional unit
$\bm{A}_{f}$  = transition Jacobian with respect to the function value $f_{t}$, nondimensional unit
$\bm{A}_{x}$  = transition Jacobian with respect to the state $\bm{x}$, nondimensional unit
$\bm{C}_{x}$  = measurement Jacobian with respect to the state $\bm{x}$, nondimensional unit
$\bar{\bm{u}}$  = augmented inducing point set, i.e., $\bar{\bm{u}}=[\bm{u}^{T},f_{t}^{T}]^{T}$, nondimensional unit
$\bm{X}_{t}$  = augmented state at time $t$, i.e., $\bm{X}_{t}=[\bm{x}_{t}^{T},{\bm{u}}^{T}]^{T}$, nondimensional unit
$\bm{\xi}_{t}$  = mean of the augmented state $\bm{X}_{t}$, nondimensional unit
$\bm{\Sigma}_{t}$  = covariance of the augmented state $\bm{X}_{t}$, nondimensional unit
$\bar{\bm{X}}_{t}$  = augmented state with $\bar{\bm{u}}$, i.e., $\bar{\bm{X}}_{t}=[\bm{x}_{t}^{T},\bar{\bm{u}}^{T}]^{T}$, nondimensional unit
$\bar{\bm{\xi}}_{t}$  = mean of the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$\bar{\bm{\Sigma}}_{t}$  = covariance of the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$\bm{\Phi}$  = transition Jacobian for the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$\bm{\Sigma}_{p,\bm{X}}$  = process noise covariance for the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$\bm{H}$  = measurement Jacobian for the augmented state $\bar{\bm{X}}_{t}$, nondimensional unit
$M$  = budget for inducing point set size, nondimensional unit
$u_{\mathrm{d}}$  = inducing point to be discarded, nondimensional unit
$\bm{u}_{\mathrm{l}}$  = inducing points left after discarding, nondimensional unit
@@
$s_{d}$  = score evaluting the importance of the inducing points, nondimensional unit
$\bm{Q}$  = inversion of the kernel matrix $\bm{K}_{uu}^{-1}$, nondimensional unit
$\bm{\Omega}$  = inversion of the joint covariance $\bm{\Sigma}_{t}^{-1}$, nondimensional unit
$\bm{Q}_{dd}$  = the $i_{\mathrm{d}}$-th diagonal element of the matrix $\bm{Q}$, nondimensional unit
$\bm{Q}_{du}$  = the $i_{\mathrm{d}}$-th row of the matrix $\bm{Q}$, nondimensional unit
$\bm{\Omega}_{dd}$  = the $(d_{x}+i_{\mathrm{d}})$-th diagonal element of the matrix $\bm{\Omega}$, nondimensional unit
$\gamma$  = novelty metric for inducing points, nondimensional unit
$\varepsilon_{\text{tol}}$  = threshold for adding inducing points, nondimensional unit
@@ 1 Introduction
State-space models (SSMs), also known as hidden Markov models (HMMs), describe the evolution and observation processes of latent states through a state transition model and a measurement model. By summarizing historical information into the latent state, SSMs provide a concise representation of system dynamics and are widely applied in fields such as aerospace engineering [1, 2], target tracking [3, 4, 5], and control systems [6, 7, 8, 9]. Nevertheless, due to the complexity of real-world systems, establishing precise SSMs based on first principles is challenging. To achieve sufficiently accurate system models, extensive research has focused on data-driven SSM modeling [10, 11, 12, 13].
@@
    Particle Filter (PF)-based method [43, 44, 45]. This approach is based on the paradigm of the Rao-Blackwellized Particle Filter [46]. It is similar to the AKF-based method, with the key difference being the use of the particle filter (PF) for the inference process. Since PF is inefficient at handling high-dimensional states, GP parameters are not augmented into the state space. Instead, each particle is assigned a GP, which is recursively learned using the particle state. This method can theoretically achieve precise inference for arbitrary observational distributions, provided the number of particles is sufficient. However, despite the inherent issues with PF, such as high computational cost and particle degeneracy [47], it also requires prior knowledge of the operating domain.
@@
To overcome the aforementioned shortcomings, a recursive learning algorithm for GPSSMs is proposed in this paper, capable of online adaptation to the operating domain and GP hyperparameters. The main contributions can be summarized as follows: Firstly, to implement a domain-independent recursive learning algorithm for GPSSMs, a novel Bayesian update equation is derived without pre-parameterizing the GP. In the derivation process, only mild linearization approximations are applied to achieve a closed-form solution, offering advantages in learning speed and stability. Through continuously expanding the inducing-point set, this method can accommodate real-time data distributions. Secondly, to address the computational cost issue caused by the increasing number of inducing points, an selection algorithm is developed with both addition and removal operations. To minimize accuracy loss, an optimal discarding equation and selection metric for inducing points are derived based on informative criteria. Thirdly, to alleviate the burden of offline hyperparameter tuning, an online GP hyperparameter optimization method is proposed. In this method, an approximate observation likelihood model is extracted from the obtained posterior distribution. Utilizing this likelihood model as the information source, the GP hyperparameters are optimized, and the posterior distribution can be adjusted as the hyperparameters change. Fourthly, to ensure numerical stability, a stable implementation method based on Cholesky decomposition is developed. Additionally, the extension to multi-output GP is provided, which significantly enhances the practicality and flexibility. Finally, the effectiveness of the proposed method is validated through two synthetic simulations and a real-world dataset. The experimental results demonstrate the superior accuracy, computational efficiency, and adaptability of the proposed method compared to state-of-the-art online GPSSM techniques.
@@ 3.2 Factorized Approximate Distribution
To address the nonlinearity within the transition and measurement models, the following theorem is derived in this work by applying first-order linearization. To the best of our knowledge, this is the first time such a result has been derived.
@@
$q(\bm{x}_{t},\bm{f})=p(\bm{f}_{\neq u}|\bm{u})q(\bm{x}_{t},\bm{u})$ (7)
where $\bm{u}\triangleq\{u_{i}\}_{i=1}^{n_{u}}$ is a collection of a finite number of inducing points, $\bm{f}_{\neq u}$ is the function values of the function $f(\cdot)$ except for $\bm{u}$, and $p(\bm{f}_{\neq u}|\bm{u})$ is the exact GP conditional distribution similar to (2), namely,
$p(\bm{f}_{\neq u}|\bm{u})=\mathcal{N}\left(\bm{K}_{f_{\neq u}u}\bm{K}_{uu}^{-1}\bm{u},\,\bm{K}_{f_{\neq u}f_{\neq u}}-\bm{K}_{f_{\neq u}u}\bm{K}_{uu}^{-1}\bm{K}_{uf_{\neq u}}\right)$ (8)
@@
Given Theorem 1, since $p(\bm{f}_{\neq u}|\bm{u})$ is known, we only need to update the finite-dimensional distribution $q(\bm{x}_{t},\bm{u})$, which build a foundation for addressing the non-parametric nature of the GP. Moreover, the approximate joint distribution remains Gaussian, which can be evaluated using closed-form update equations. In the following, we will prove Theorem 1 and then derive these equations.
@@
[…]
@@
By means of the above notation, we will illustrate the proof of Theorem 1, which is based on mathematical induction. First, it can be demonstrated that at time 0, the joint distribution naturally satisfies the factorized form (7):
@@
where we use the notation for moments in (9) and the shorthand notation $f_{t}=f(\bm{\mu}_{t})$, $m_{f_{t}}=\mathbb{E}_{q(\bm{x}_{t},\bm{f})}[f_{t}]$, $\bm{F}_{t}=\bm{F}(\bm{\mu}_{t},m_{f_{t}})$. Additionally, the two Jacobian matrices in the equation are
@@ 3.3 Moment Matching Equation
where $\bar{\bm{\xi}}_{t+1}^{-}$ and $\bar{\bm{\Sigma}}_{t+1}^{-}$ represent the predicted mean and covariance for the augmented state $\bar{\bm{X}}_{t+1}$. $\bm{\Phi}$ denotes the transition Jacobian matrix for $\bar{\bm{X}}_{t+1}$, $\bm{\Sigma}_{f,\bar{X}}$ is the process noise covariance for $\bar{\bm{X}}_{t+1}$, and $\bm{I}_{n_{u}}$ is the identity matrix.
@@ 4.1 Dynamic Adjusting the Inducing Points Set
$\hat{q}(\bm{x},\bm{f})=p(\bm{f}_{\neq u_{\mathrm{l}}}|\bm{u}_{\mathrm{l}})\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})$ (19)
[…]
@@
$=\arg\min\limits_{\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})}\mathrm{KL}\left[p(\bm{f}_{\neq u}|\bm{u})q(\bm{x},\bm{u})\|p(\bm{f}_{\neq u_{\mathrm{l}}}|\bm{u}_{\mathrm{l}})\hat{q}(\bm{x},\bm{u}_{\mathrm{l}})\right]$
@@ 4.2 Online Optimization of GP Hyperparameters
where $q_{\text{new}}(\bm{x}_{t},\bm{f})$ denotes the approximate joint distribution corresponding to the new hyperparameters $\bm{\theta}_{\text{new}}$. In (28), since the distributions $p(\bm{f};\bm{\theta}_{\text{old}})$, $p(\bm{f};\bm{\theta}_{\text{new}})$, and $q_{\text{old}}(\bm{x}_{t},\bm{f})$ are all Gaussian, the new approximate distribution $q_{\text{new}}(\bm{x}_{t},\bm{f})$ is also Gaussian. By expanding the distribution related to $\bm{f}$ in the third row of (28) with the conditional distribution $p(\bm{f}|\bm{u})$, it can be shown that the new distribution retains the factorized form $q_{\text{new}}(\bm{x}_{t},\bm{f})=p(\bm{f}_{\neq u}|\bm{u};\bm{\theta}_{\text{new}})q_{\text{new}}(\bm{x}_{t},\bm{u})$, where: